OpenStudy (anonymous):

PLEASE HELP! (ALL MULTIPLE CHOICE) Which equation represents a line perpendicular to the graph of the line -2y=4x+10 A) y=x+6 B) y=-1/2x+6 C)y=2x+6 D)y=-2x+6 The expression (x-4)^2 is equivalent to A)x^2-16 B) x^2+16 C) x^2-8x+16 D) x^2+8x+16

OpenStudy (anonymous):

ALSO THIS ONE PLEASE!

OpenStudy (anonymous):

IT IS MULTIPLE CHOICE QUESTION (PDF FILE BECAUSE IT HAS A GRAPH)

OpenStudy (anonymous):

The answer to the second question is C

OpenStudy (anonymous):

i thought it was A

jimthompson5910 (jim_thompson5910):

start off by solving -2y=4x+10 for y -2y=4x+10 y=4x/(-2)+10/(-2) y = -2x - 5

jimthompson5910 (jim_thompson5910):

the slope of y = -2x - 5 is -2 anything perpendicular to this will have a slope of 1/2 (flip the fraction and flip the sign)

OpenStudy (anonymous):

so the answer would be A?

jimthompson5910 (jim_thompson5910):

assuming A is a typo and it should have 1/2x, then yes

OpenStudy (anonymous):

so it would not be -1/2x+6? It was not a typo.

OpenStudy (anonymous):

and can u help me with the attached file in the comments? It is an absolute value

jimthompson5910 (jim_thompson5910):

the slope of the perpendicular line is positive (since the original line has a negative slope)

OpenStudy (anonymous):

i see i will ask my teacher can u @jim_thompson5910 help me with the absolute value screenshot in the comments. it is a multiple choice graph

OpenStudy (anonymous):

 The arrows mean you multiply those things